OpenStudy (anonymous):

Can someone please graph this?? Find the inverse of the function below. Graph the function below and the inverse function. Determine the domain, range, and asymptotes of the function below and the inverse function. Please show your work. f(x) = 3 ln(4-x)+1

OpenStudy (anonymous):

Here's the answer: Is this correct?!! y = 3 ln(4-x)+1, solving for x (y-1)/3 = ln(4-x) taking exponential power on both sides we get e^[(y-1)/3] = e^ln(4-x) e^[(y-1)/3] = (4-x) x=4-e^[(y-1)/3] interchange x to y y =4-e^[(x-1)/3] answer....inverse function from the function f(x) = 3 ln(4-x)+1 vertical retriceymptote x=4 range(-∞,∞)

OpenStudy (turingtest):

everything look good. here's a graph http://www.wolframalpha.com/input/?i=+f%28x%29+%3D+3+ln%284-x%29%2B1 If you don't already know, WolframAlpha is a great way to check answers

OpenStudy (fretje):

I can give some couples of the function:(-10,8.9171),(-5,7.5916),(0,5.1589),(2,3.0794),(3,1),(3.5,-1.0794),(3.99,-12.8155)

OpenStudy (anonymous):

Thank you!!

OpenStudy (fretje):

and the inverse function f-1(y) is correct and for the function f(x) asymptote is x=4 and domain is thus x<4 and the range is -Infinity to Infinity so what you had was correct unless you want the complex numbers where x>4 is allowed.

OpenStudy (fretje):

for the inverse function I have to ponder some more

OpenStudy (turingtest):

She already has the answer fretje, she got it right.